OpenStudy (anonymous):

Why is \[cos(arctan(x))=\frac{1}{\sqrt{x^2+1} } \] ?

Parth (parthkohli):

Let \(\arctan(x)\) be \(\rm y\).\[\rm \cos (y)\]Can you find an expression for \(\rm cos(y)\) in terms of \(\rm sine\)?

OpenStudy (anonymous):

cosy=sin(y+pi/2)

Parth (parthkohli):

I'd do this instead...\[\rm \sqrt{1 - \sin^2(y)}\]

OpenStudy (anonymous):

OK

Parth (parthkohli):

I'm stuck on this too. :p

OpenStudy (anonymous):

\[1-s^2arctanx=1/(x^2+1)\] No, it seems too messy

OpenStudy (turingtest):

you must be working with this triangle

OpenStudy (anonymous):

Not knowingly, but I suppose so.

OpenStudy (anonymous):

Genius solution! Thanks.

OpenStudy (turingtest):

\[\tan\theta=\frac x1=x\implies\theta=\tan^{-1}x\]\[\cos\theta=\frac1{\sqrt{x^2+1}}\implies\cos(\tan^{-1}x)=\frac1{\sqrt{x^2+1}}\]not sure where this triangle is coming from welcome!

OpenStudy (anonymous):

How often is it very useful (i.e. hard not to solve while not) thinking geometrically with trig identities? That is, is it worth doing often or only when I see sqrt(1+x^2) lying about?

OpenStudy (turingtest):

I guess you did an integration and substituted x=tan(u) or something. Whenever I do trig substitutions I *always* draw the triangle to reduce things into simpler terms of x at the end, but I don't memorize anything except the trig definitions ant the Pythagorean theorem for that.

OpenStudy (turingtest):

and*

OpenStudy (anonymous):

OK- it was in fact Taylor seriesing for arctan. I'm not sure where the cosine popped up, though